OpenStudy (anonymous):

what is the distance between point R at (6,3) and y=2x-6 ??? show work please

OpenStudy (anonymous):

the closest y=2x+6 comes to 6,3 is 5,4... But the function is not one point...

OpenStudy (anonymous):

but i think the distance is \[\sqrt{2}\]

OpenStudy (anonymous):

k to find the distance think of the line connecting (6,3) with the line y = 2x - 6. it will be perpendicular to that line, because "distance" means minimum distance.

OpenStudy (anonymous):

the equation for that line will have slope \[-\frac{1}{2}\] because it is perpendicular. and it will contain the point (6,3)

OpenStudy (anonymous):

so the equation will be \[y-3=-\frac{1}{2}(x-6)\] \[y-3=-\frac{1}{2}x + 3\] \[y = -\frac{1}{2}x+6\]

OpenStudy (anonymous):

now we can find the point of intersection by setting the two lines equal and solving for x: \[2x-6=-\frac{1}{2}x+6\] \[4x-12=-x+12\] \[5x=24\] \[x=\frac{24}{5}\]

OpenStudy (anonymous):

if you are still there you should check my algebra because i often make mistakes.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

of course we are still not done. we know x, now find y and you will have the point on the line closest to (6,3) then you have to use the distance formula

OpenStudy (anonymous):

ok so lets give dat a shot

OpenStudy (anonymous):

hold on a second there may be something wrongn with my algebra

OpenStudy (anonymous):

easier if i write it on paper. the line was y = 2x - 6 and the point is (6,3 )yes?

OpenStudy (anonymous):

no it is good. the numbers coordinates are (4.8,3.6)

OpenStudy (anonymous):

now distance formula yes?

OpenStudy (anonymous):

yeah im ready sorry for takin long to replie i couldnt figuer out how to get back 2 dis page

OpenStudy (anonymous):

1.34164 if we are working with decimals

OpenStudy (anonymous):

no decimals in this

OpenStudy (anonymous):

ok then i guess we work with the fractions

OpenStudy (anonymous):

\[\sqrt{(6-\frac{24}{5})^2+(3-\frac{18}{5})^2}\]

OpenStudy (anonymous):

how about i scan the work sheet on here u ok with that ill show u da question

OpenStudy (anonymous):

\[\sqrt{(\frac{6}{5})^2+(\frac{3}{5})^2}\]

OpenStudy (anonymous):

\[\sqrt{\frac{45}{25}}\]

OpenStudy (anonymous):

\[\frac{3\sqrt{5}}{5}\]

OpenStudy (anonymous):

sure but i like my answer

OpenStudy (anonymous):

cool but jus look it look close

OpenStudy (anonymous):

great! the answer is there and it is D because \[\frac{3\sqrt{5}}{5}=\frac{3}{\sqrt{5}}\]

OpenStudy (anonymous):

thank u so much!!!!